As for Ranchi, it has a population of 45L as per the last census taken in 2010. The year is 2020, and the company is looking to open a new footwear production factory. In Ranchi, Foot&Bros and Slipper&Sons are the current market leaders. Foot&Bros have been in business for 40+ years and have a 60% market share. In Ranchi, Slipper&Sons hold a 40% market share. Excess Industries expects to capture 25% of the former and 40% of the latter.
Factory lease (99 years) 99 Cr. Rent 1.5 L/month Electricity 10 units/footwear produced Electricity rate is 15 rupees per unit in Ranchi.
We will be sourcing our raw materials from two locations - nearby Kolkata (50%) and Patna (50%) . We understand that each shoe requires 10m of leather for manufacturing reasons.Wastage will be 10m per shoe.
Profit margin 20% - Analyze and estimate the profitability, revenues and costs of Excess industries from the beginning of 2021 until 2021 (first four years of business).
1
Expert's answer
2021-09-04T15:16:44-0400

Given

"A_1=50000\\N=25\\i=0.07\\g=0.05"

therefore

"P=A_1[\frac{1-(1+g)^N(1+I)^{-N}}{i-g}]\\P=50000[\frac{1-(1+0.05)^{25}(1+0.07)^{-25}}{0.07-0.05}]\\=940167.22"

Benefit paid after 25 years

"A=50000\\N=25\\i=0.07"

Present value of retirement benefits

"=50000(P/A,7\%,25)\\=50000(11.654)\\=582700"


Additional saving"=940167.22-582700=357467"